OpenStudy (anonymous):

Tell wether the ratios are equivalent or not equivalent 16:3,27:5

OpenStudy (texaschic101):

ratios are nothing but fractions 16:3 is the same as 16/3 27:5 is the same as 27/5 to see if they are equivalent, lets make a proportion 16/3 = 27/5 --- cross multiply (27)(3) = (16)(5) 81 = 80 (incorrect) --- these are not equivalent